| 1. Let z = 1+ 2i z = -2-2i, z = 3, 24 =i A. Complex arithmetic (20%) | a. Zi + Z2 b. Z1Zz sle Isles
B. Determine the principle value of the argument and graph it (20%) a. 21 b. Z2 c. 23 d. 24
Answer #1

-1 1. A. () 2+27 = 1+21+ (-2-21) = (1-2) +i (2-2) = -171.0 6) 3772 = (1+2i)(-21-21) 1,= (1+z) (-212i) - (-21) (-2+2i) -2 +20